OpenStudy (anonymous):

Solve the following quadratic equation. 3(x + 5)^2 = 60

hartnn (hartnn):

start by dividing both sides by 3 what u get ?

OpenStudy (anonymous):

Well 60/3 is 20... how do you divide the other side by 3?

hartnn (hartnn):

3*something / 3 = something

OpenStudy (anonymous):

1(x+5)^2?

hartnn (hartnn):

yes so you now have \(\Large (x+5)^2 =20\) take square root on both sides, what u get ?

OpenStudy (anonymous):

(x+2) ^2=4?

hartnn (hartnn):

where did that come from ? :O

hartnn (hartnn):

let me give you a hint : \(\Large if, a^2 =b \\ \Large then, a= b \quad or \quad a=-b \)

hartnn (hartnn):

if \((x+5)^2 =20\) then ???